### v2

# A generalisation of the Euclid–Mullin sequences

Abstract.
We extend Mullin’s prime-generating procedures to produce sequences of primes lying in given residue classes. In particular we study the sequences generated by cyclotomic polynomials $\Phi_{m}(cx)$ for suitable $c\in\mathbb{Z}$. Under the Extended Riemann Hypothesis in general and unconditionally for some moduli, we show that the analogue of the second Euclid–Mullin sequence omits infinitely many primes $\equiv 1\pmod{m}$. We further show unconditionally that at least one prime is omitted for infinitely many $m$. This generalises work of the first author for $m=1$ and the second author for $m=2^{k}$.
1. Introduction
In [10], Mullin constructs two sequences of prime numbers based on Euclid’s proof of infinitude of primes. For Mullin’s first sequence, say $\{p_{n}\}_{n=1}^{\infty}$, define $p_{n}$ to be the smallest prime dividing $p_{1}\cdots p_{n-1}+1$ (with the convention that the empty product is $1$, so $p_{1}=2$). Mullin’s second sequence, $\{P_{n}\}_{n=1}^{\infty}$, is defined analogously, choosing instead the largest prime at each stage. Mullin then asked whether every prime is contained in each of these sequences, and if not, whether they are recursive, i.e. whether there is an algorithm to determine membership of a given prime.
Cox and van der Poorten [6] showed that the second sequence omits all primes less than $47$, apart from the first four terms $2$, 33, 77, and 4343.111Their paper claimed to show that $47$ is omitted as well, but contained a numerical error that went unnoticed for decades; we thank Thorkil Naur for bringing this to our attention. With the terms of the sequence that were known at the time, the Cox–van der Poorten method only works to rule out primes below $47$, and it only became possible to rule $47$ out with the computation of the $12$th term by Wagstaff [22] in 1991. Their method can be extended to the primes less than $79$ using more known terms of $\{P_{n}\}_{n=1}^{\infty}$ [7, 13, 22, 3]. Later, the first author [4] showed that the second sequence omits infinitely many primes, and Pollack and Treviño [15] provided a proof relying only on elementary number theory. Mullin’s second question remains open, but the first author showed in [4, Theorem 2] that if $\{P_{n}\}_{n=1}^{\infty}$ is not recursive then it must have Dirichlet density zero in the primes.
As for the first sequence, both of Mullin’s questions remain open, and very little is known. Shanks [17] conjectured that the first sequence does indeed contain every prime, as otherwise the cumulative products $p_{1}\cdots p_{n}$ would always be invertible modulo the smallest omitted prime $q$, but would conspire to avoid the residue class $1\pmod{q}$ for all but finitely many $n$. This observation formed the inspiration for this paper, as we now consider what happens should we only generate primes in a fixed residue class.
A general framework for studying Euclidean-style proofs of the infinitude of primes in residue classes was described by Pollack [14], generalising earlier work of Murty [12] (see also [11]). Given a coprime progression $a\bmod m$, Pollack calls a polynomial $f\in\mathbb{Z}[x]$ an $E^{\prime}(a,m)$ polynomial if
  • (i)
    $f(n)$ has a prime divisor $p\equiv a\pmod{m}$ for every sufficiently large $n\in\mathbb{N}$;
  • (ii)
    for every $p\equiv a\pmod{m}$, there exists $n\in\mathbb{Z}$ such that $p\nmid f(n)$.
(We will also write $f\in E^{\prime}(a,m)$ to mean the same.) Pollack proceeded to show that any $E^{\prime}(a,m)$ polynomial yields a Euclidean-style proof of infinitude of primes $p\equiv a\pmod{m}$. From work of Schur [16], we know that $E^{\prime}(a,m)$ polynomials exist when $a^{2}\equiv 1\pmod{m}$, and Pollack proved that the converse holds assuming Schinzel’s Hypothesis H.
Pollack’s proof of infinitude of primes $p\equiv a\pmod{m}$ given $f\in E^{\prime}(a,m)$ involves an arbitrary choice at each step of the construction, and is therefore not well suited to our goal of producing explicit, uniquely-defined sequences. For that purpose, we make yet another definition:
Definition 1.1.
We say that $f\in\mathbb{Z}[x]$ is a Generalised Euclid–Mullin polynomial for the progression $a\bmod m$, and write $f\in\mathrm{GEM}(a,m)$, if
  • (i)
    $f(1)\neq 0$ and
  • (ii)
    for every $n\in\mathbb{N}$ satisfying $(n,f(0))=1$ and $n\equiv a^{k}\pmod{m}$ for some $k\in\mathbb{N}$, $f(n)$ has a prime divisor $p$ satisfying $p\nmid f(0)$ and $p\equiv a\pmod{m}$.
Given $f\in\mathrm{GEM}(a,m)$, we can construct a sequence of primes $(p_{n})_{n\geq 1}$ by choosing $p_{n}$ to be a prime divisor of $f(p_{1}\cdots p_{n-1})$ satisfying $p_{n}\nmid f(0)$ and $p_{n}\equiv a\pmod{m}$. Since $f(1)\neq 0$ and
$f(p_{1}\cdots p_{n-1})\equiv f(0)\pmod{p_{1}\cdots p_{n-1}}$,
it is clear that $f(p_{1}\cdots p_{n-1})$ never vanishes, and that the $p_{n}$ are distinct, since $p_{n}\nmid f(0)\implies p_{n}\nmid p_{1}\cdots p_{n-1}$.
We remark that Definition 1.1 is not the most general one that would enable the construction of infinite sequences of distinct primes $p\equiv a\pmod{m}$, but it has the added feature that we can prove a sort of equivalence between $\mathrm{GEM}(a,m)$ and $E^{\prime}(a,m)$. More precisely, although they are not the same as sets, they generate the same proofs, in the sense that an $E^{\prime}$ polynomial can be converted into a $\mathrm{GEM}$ polynomial, and vice versa:
Proposition 1.2.
For any coprime progression $a\bmod m$,
$E^{\prime}(a,m)\neq\emptyset\iff\mathrm{GEM}(a,m)\neq\emptyset.$
Proof.
Let $f\in E^{\prime}(a,m)$. Choose $c\in\mathbb{Z}$ such that $f(c)\neq 0$, and set
$\ell=\operatorname{lcm}\{p-1:p\mid f(c),\;p\equiv a\pmod{m}\}$.
By hypothesis, for each $p\mid f(c)$ with $p\equiv a\pmod{m}$, there exists $n_{p}\in\mathbb{Z}$ with $p\nmid f(n_{p})$. By the Chinese Remainder Theorem, we can choose a single $A\in\mathbb{N}$ such that $A+c\equiv n_{p}\pmod{p}$ for every such $p$. Furthermore, we may take $A$ sufficiently large to ensure that $f(n)$ is nonzero and has a prime factor $\equiv a\pmod{m}$ for every $n\geq A+c$.
Now consider $F(x)=f(Ax^{\ell}+c)$, and take $n\in\mathbb{N}$ coprime to $F(0)=f(c)$. Then $F(n)=f(An^{\ell}+c)$ is nonzero and has a prime divisor $p\equiv a\pmod{m}$. If $p\mid F(0)$ then $n^{\ell}\equiv 1\pmod{p}$ by Fermat’s Little Theorem, and therefore $F(n)\equiv f(A+c)\equiv f(n_{p})\not\equiv 0\pmod{p}$. This is a contradiction, so we must have $p\nmid F(0)$. Since $n$ was arbitrary, $F\in\mathrm{GEM}(a,m)$.
Conversely, suppose $F\in\mathrm{GEM}(a,m)$, and set
$C=|F(0)|,\quad D=(F(1),C^{\infty})=\prod_{p\mid C}p^{v_{p}(F(1))},\quad f(x)=D^{-1}F(1+CDmx).$
Considering the Taylor expansion of $F$ around $1$, we see that $f$ has integral coefficients. Let $p$ be a prime number with $p\equiv a\pmod{m}$. If $p\mid F(0)$ then by the definition of $D$, $p\nmid D^{-1}F(1)=f(0)$. If $p\nmid F(0)$ then choosing $n_{p}\in\mathbb{Z}$ to satisfy $1+CDmn_{p}\equiv 0\pmod{p}$ (which we can do since $p\nmid CDm$), we see that $p\nmid f(n_{p})$.
Now let $n\in\mathbb{N}$. Then by hypothesis $F(1+CDmn)$ has a prime divisor $p\equiv a\pmod{m}$ such that $p\nmid C$. Since all prime divisors of $D$ divide $C$, $p$ is not among them, and thus $p\mid f(n)$. Therefore $f\in E^{\prime}(a,m)$. ∎
Two natural sequences are obtained from Mullin’s decree of choosing $p_{n}$ as small as possible at each step, or as large as possible. We call these the first and second generalised Euclid–Mullin sequences associated to $f$, and denote them $\mathrm{GEM}_{1}(f;a,m)$ and $\mathrm{GEM}_{2}(f;a,m)$. In light of Schur’s result and the known and conjectured properties of the original Euclid–Mullin sequences, we conjecture the following:
Conjecture 1.3 (Generalised Euclid–Mullin conjecture).
Let $a,m\in\mathbb{N}$ with $a^{2}\equiv 1\pmod{m}$. Then
  • (i)
    there exists $f\in\mathrm{GEM}(a,m)$ such that $\mathrm{GEM}_{1}(f;a,m)$ contains every prime $p\equiv a\pmod{m}$;
  • (ii)
    for any $f\in\mathrm{GEM}(a,m)$, $\mathrm{GEM}_{2}(f;a,m)$ omits infinitely many primes $p\equiv a\pmod{m}$.
In this paper we make progress on the second of these conjectures. Our first result is a partly conditional proof of Conjecture 1.3(ii) for $m\in\{1,2\}$. Recall that the Extended Riemann Hypothesis (ERH) is the statement that for any number field $L$, the Dedekind zeta function $\zeta_{L}(s)$ does not vanish for $\Re(s)>\frac{1}{2}$.
Theorem 1.4.
Let $m\in\{1,2\}$ and $f\in\mathrm{GEM}(1,m)$. Assume either that ERH is true or that $f$ is not of the form $c(ax+b^{8})^{k}$ with $c\in\mathbb{Z}$, $a,b,k\in\mathbb{N}$, $(a,b)=1$, and $a$ odd or divisible by $32$. Then $\mathrm{GEM}_{2}(f;1,m)$ omits infinitely many primes.
Next we study the sequences generated by the cyclotomic polynomials $\Phi_{m}(cx)$ for suitable $c\in\mathbb{Z}$. This generalises the second author’s master’s thesis [19], which treated the case when $m$ is a power of $2$. A common choice in proofs of the infinitude of primes $\equiv 1\pmod{m}$ is $\Phi_{m}(mx)$, which is a $\mathrm{GEM}(1,m)$ polynomial for all $m>1$. (For $m=1$, the original Euclid–Mullin sequences are generated from $\Phi_{1}(-x)$.) In fact for any $c$ divisible by the largest prime factor of $m$, every prime $p$ dividing $\Phi_{m}(cn)$ for some $n$ satisfies $p\equiv 1\pmod{m}$. However, there are many more possible choices, as the following proposition shows.
Proposition 1.5.
Let $m\in\mathbb{N}$ and $c\in\mathbb{Z}$. Then $\Phi_{m}(cx)$ is a $\mathrm{GEM}(1,m)$ polynomial if and only if $c\in S(m)$, where
$S(m)\coloneq\begin{cases}\mathbb{Z}\setminus\{0,1,2\}&\text{if }m=1,\\ 2\mathbb{N}\cup\bigl\{c\in-\mathbb{N}:c\nmid 2^{k}+1\text{ for all }k\geq 0\bigr\}&\text{if }m=2,\\ \mathbb{Z}\setminus\{-2,-1,0,1\}&\text{if }m=3,\\ \mathbb{Z}\setminus\{-1,0,1,2\}&\text{if }m=6,\\ \mathbb{Z}\setminus\{-1,0,1\}&\text{otherwise}.\end{cases}$
Proof.
Combining Zsigmondy’s theorem [24] with the identities
$m\equiv 1\pmod{2}$ $\implies\Phi_{m}(-x)=\pm\Phi_{2m}(x),$
$m\equiv 2\pmod{4}$ $\implies\Phi_{m}(-x)=\pm\Phi_{m/2}(x),$
$m\equiv 0\pmod{4}$ $\implies\Phi_{m}(-x)=\Phi_{m}(x),$
we see that for any $x\in\mathbb{Z}\setminus\{-1,0,1\}$, $\Phi_{m}(x)$ is nonzero and has a prime divisor $p\equiv 1\pmod{m}$ except in the following cases:
  • $m=1$ and $x=2$;
  • $m=2$ and $|x+1|$ is a power of $2$;
  • $m=3$ and x=2x=-2;
  • $m=6$ and $x=2$.
One can also see that the values of $\Phi_{m}(x)$ for $x\in\{-1,0,1\}$ are never suitable except when (m,x)=(1,1)(m,x)=(1,-1).
In order for $\Phi_{m}(cx)$ to be $\mathrm{GEM}(1,m)$ polynomial, we need $cn$ to avoid the exceptions above for every natural number $n\equiv 1\pmod{m}$. For $m\neq 2$ it is straightforward to see that this is equivalent to $c\in S(m)$.
For $m=2$ and odd $c>0$, there are infinitely many odd $n\in\mathbb{N}$ such that $cn+1$ is a power of $2$ (choose $k\in\mathbb{N}$ divisible by the order of $2\bmod c$ and set $n=(2^{k}-1)/c$). Similarly, for $m=2$ and $c<0$, there exists odd $n\in\mathbb{N}$ such that $|cn+1|$ is a power of $2$ if and only if $c\mid 2^{k}+1$ for some $k$. ∎
Again we can give a conditional proof of Conjecture 1.3(ii) in this setting:
Theorem 1.6.
Let $m\in\mathbb{N}$ and $c\in S(m)$. Then, assuming ERH, $\mathrm{GEM}_{2}(\Phi_{m}(cx);1,m)$ omits infinitely many primes $p\equiv 1\pmod{m}$.
For $m>2$ we will see that we can reduce to the Generalised Lindelöf Hypothesis (GLH) for Dirichlet $L$-functions, which is implied by ERH. Moreover, we can remove all conditional hypotheses in some cases:
Theorem 1.7.
Let $m\in\{1,\ldots,10,12,14,18\}$ and $c\in S(m)$, with the following further assumptions:
$m=1$ $\implies c\in\mathbb{N}\cup(2\mathbb{Z}\setminus 32\mathbb{Z})\cup\{-5\}\cup\bigl\{1-2^{k}:k\in\mathbb{N}\bigr\},$
$m=2$ $\implies c\in-\mathbb{N}\cup(2\mathbb{Z}\setminus 32\mathbb{Z}).$
Then $\mathrm{GEM}_{2}(\Phi_{m}(cx);1,m)$ omits infinitely many primes $p\equiv 1\pmod{m}$.
(The restrictions on $c$ for $m\in\{1,2\}$ arise because $\Phi_{m}(cx)$ has degree $1$; in view of Theorem 1.4, that is the only obstruction to giving an unconditional proof of the conjecture in those cases.)
Finally, we show unconditionally that at least one prime is omitted for infinitely many $m$:
Theorem 1.8.
Let $m>2$ be a power of $2$, and let $c\in S(m)$. Then $\mathrm{GEM}_{2}(\Phi_{m}(cx);1,m)$ omits the smallest prime $p\equiv 1\pmod{m}$ if $(m,c)\notin\{(4,\pm 2),(4,\pm 3),(8,\pm 2)\}$, and the second smallest such prime otherwise.
1.1. Outline
The contents of the paper are as follows:
  • In Section 2 we develop the character sum machinery necessary for Theorem 1.7, culminating in the proof of Theorem 2.5. This is modelled on the proof of [4, Lemma 4], and the main ingredient is again Burgess’ estimate [5] for short character sums, leveraged by the Dirichlet hyperbola method.
  • Section 3 derives a similar result for higher-order characters, Theorem 3.2, to be used in the proof of Theorems 1.4 and 1.6. Since progress toward GLH for Dedekind zeta functions of non-abelian extensions of $\mathbb{Q}$ is much more limited than for abelian extensions, we derive this conditionally on ERH.
  • Section 4 derives or quotes a few miscellaneous results from Diophantine equations that are needed in the proofs of Theorems 1.41.8.
  • Finally, we assemble these ingredients in Section 5 to complete the proofs.
Acknowledgements
The first author is grateful to Dan Fretwell for suggesting the study of cyclotomic analogues of the Euclid–Mullin sequences.
2. Unconditional analytic estimates
The goal of this section is to prove an analogue of [4, Lemma 4] adapted to the setting of cyclotomic polynomials; see Theorem 2.5. Fix $m\in\mathbb{N}$, let $K=\mathbb{Q}(e^{2\pi i/m})$ be the $m$th cyclotomic field, and denote its ring of integers by $\mathcal{O}_{K}=\mathbb{Z}[e^{2\pi i/m}]$. Let $\zeta_{K}(s)$ denote the Dedekind zeta function of $K$, so that
$\zeta_{K}(s)=\prod_{\mathfrak{p}}\frac{1}{1-N(\mathfrak{p})^{-s}},$
where the product runs over prime ideals $\mathfrak{p}\subset\mathcal{O}_{K}$. Define coefficients $a(n)$ by
$\zeta_{K}(s)=\sum_{n=1}^{\infty}\frac{a(n)}{n^{s}}.$
Further set $T=\{\mathfrak{p}:\mathfrak{p}\mid m\mathcal{O}_{K}\text{ and }f(\mathfrak{p})=1\}\cup\{\mathfrak{p}:f(\mathfrak{p})\text{ is odd and}>1\}$, where $f(\mathfrak{p})$ denotes the inertia degree of $\mathfrak{p}$, and define $a^{\flat}(n)$ such that
$\zeta_{K}(s)\prod_{\mathfrak{p}\in T}(1-N(\mathfrak{p})^{-s})=\sum_{n=1}^{\infty}\frac{a^{\flat}(n)}{n^{s}}.$
Lemma 2.1.
Let $\chi$ be a character of modulus $q$ and conductor $q^{\ast}>1$. Let $q_{1}$ be the maximal cubefree unitary divisor of $q^{\ast}$, and set $q_{2}=q^{\ast}/q_{1}$, $q_{0}=\prod_{\begin{subarray}{c}p\mid q\\ p\nmid q^{\ast}\end{subarray}}p$. Then
$\sum_{n=M+1}^{M+N}\chi(n)\ll_{r,\varepsilon}q_{0}^{\varepsilon}q_{1}^{\frac{r+1}{4r^{2}}+\varepsilon}q_{2}^{\frac{1}{r}}N^{1-\frac{1}{r}}\quad\text{for all }M,N,r\in\mathbb{N}\text{ and }\varepsilon>0.$
Proof.
When $\chi$ is primitive and $q_{2}=1$, this is the result of Burgess [5, Theorem 2]. As explained in [8, (12.56)], we can extend to non-principal primitive characters by breaking into progressions modulo $q_{2}$ and applying Burgess’ result. Finally, we can extend to all non-principal characters by sieving:
$\sum_{n=M+1}^{M+N}\chi(n)=\sum_{n=M+1}^{M+N}\sum_{d\mid(n,q_{0})}\mu(d)\chi^{\ast}(n)=\sum_{d\mid q_{0}}\mu(d)\chi^{\ast}(d)\sum_{\frac{M}{d}<n\leq\frac{M+N}{d}}\chi^{\ast}(n)\ll_{r,\varepsilon}q_{0}^{\varepsilon}q_{1}^{\frac{r+1}{4r^{2}}+\varepsilon}q_{2}^{\frac{1}{r}}N^{1-\frac{1}{r}}.$
Lemma 2.2.
Let $\chi$ be a non-trivial character of modulus $q$. Assuming GLH for Dirichlet $L$-functions, we have
$\sum_{n\leq x}\chi(n)\ll_{\varepsilon}q^{\varepsilon}\sqrt{x}.$
Proof.
By periodicity we may assume without loss of generality that $x\leq q$. By Perron’s formula [21, Corollary 2.1], for $x\geq 2$,
$\sum_{n\leq x}\chi(n)=\frac{1}{2\pi i}\int_{1+\frac{1}{\log{x}}-ix}^{1+\frac{1}{\log{x}}+ix}L(s,\chi)x^{s}\,\frac{ds}{s}+O(\log{x}).$
In turn, under GLH we have $L(s,\chi)\ll_{\varepsilon}|qs|^{\frac{\varepsilon}{2}}$ uniformly for $\Re(s)\geq\frac{1}{2}$. Shifting the contour to $\Re(s)=\frac{1}{2}$, we see that
$\sum_{n\leq x}\chi(n)\ll_{\varepsilon}(qx)^{\frac{\varepsilon}{2}}\sqrt{x}\leq q^{\varepsilon}\sqrt{x}.$
Lemma 2.3.
Let $\chi$ be a quadratic character modulo $q$, not necessarily primitive, with conductor not dividing $m$. For $r\in\mathbb{N}$, define
$\nu(r)=\begin{cases}0&\text{if $r=2$ and GLH holds for Dirichlet $L$-functions},\\ \frac{r+1}{4r^{2}}&\text{otherwise}.\end{cases}$
Then for any $r\in\mathbb{N}$ and $\varepsilon>0$,
$\sum_{n\leq x}a^{\flat}(n)\chi(n)\ll_{m,r,\varepsilon}q^{\nu(r)+\varepsilon}x^{1-\frac{1}{\varphi(m)r}}.$
Proof.
We first show that it suffices to prove the corresponding estimate with $a^{\flat}(n)$ replaced by $a(n)$. To that end, define coefficients $b(n)$ such that
$\prod_{\mathfrak{p}\in T}(1-N(\mathfrak{p})^{-s})=\sum_{n=1}^{\infty}\frac{b(n)}{n^{s}},$
and note that
$\sum_{n=1}^{\infty}\frac{|b(n)|}{n^{\sigma}}<\infty\quad\text{for all }\sigma>\tfrac{1}{3}.$
Then $a^{\flat}=a\ast b$, so assuming the desired result holds for $a(n)$, we have
$\sum_{n\leq x}a^{\flat}(n)\chi(n)$ $=\sum_{n\leq x}\sum_{d\mid n}b(d)a(n/d)\chi(n)=\sum_{d\leq x}b(d)\chi(d)\sum_{n\leq\frac{x}{d}}a(n)\chi(n)$
$\ll_{m,r,\varepsilon}\sum_{d\leq x}|b(d)\chi(d)|q^{\nu(r)+\varepsilon}\left(\frac{x}{d}\right)^{1-\frac{1}{\varphi(m)r}}\leq q^{\nu(r)+\varepsilon}x^{1-\frac{1}{\varphi(m)r}}\sum_{d=1}^{\infty}\frac{|b(d)|}{d^{1-\frac{1}{\varphi(m)r}}}.$
The final series over $d$ always converges, since $1-\frac{1}{\varphi(m)r}>\frac{1}{3}$ unless $m\leq 2$, and in those cases there are no primes $\mathfrak{p}$ with $f(\mathfrak{p})>1$, so $b(d)=0$ for $d>2$.
Let $\xi_{1},\ldots,\xi_{\varphi(m)}$ denote the primitive characters of conductor dividing $m$, and put $\chi_{i}=\xi_{i}\chi$, so that $a(n)\chi(n)=(\chi_{1}\ast\cdots\ast\chi_{\varphi(m)})(n)$. We will prove by induction on $k\geq 1$ that
(2.1) $\sum_{n\leq x}(\chi_{1}\ast\cdots\ast\chi_{k})(n)\ll_{k,m,r,\varepsilon}q^{\nu(r)+\varepsilon}x^{1-\frac{1}{kr}}.$
The result then follows on taking $k=\varphi(m)$.
For a single character $\chi_{i}=\xi_{i}\chi$, first note that since $\chi$ is quadratic, its conductor is cubefree apart from a possible factor of $8$. Thus, if $q_{0},q_{1},q_{2}$ are as in Lemma 2.1 applied to $\chi_{i}$ then $q_{2}\mid 8m$. Hence, by Lemmas 2.1 and 2.2, we have
$\sum_{n\leq x}\chi_{i}(n)\ll_{m,r,\varepsilon}q^{\nu(r)+\varepsilon}x^{1-\frac{1}{r}},$
which implies the $k=1$ case of (2.1).
Suppose (2.1) holds for some $k<\varphi(m)$, and write
$f=\chi_{1}\ast\cdots\ast\chi_{k},\quad g=\chi_{k+1},\quad F(x)=\sum_{n\leq x}f(n),\quad G(x)=\sum_{n\leq x}g(n).$
Then by the Dirichlet hyperbola method, for any $y>0$, we have
$\sum_{n\leq x}(f\ast g)(n)=\sum_{n\leq y}F\!\left(\frac{x}{n}\right)g(n)+\sum_{n\leq\frac{x}{y}}f(n)G\!\left(\frac{x}{b}\right)-F\!\left(\frac{x}{y}\right)G(y).$
Taking $y=x^{\frac{1}{k+1}}$ and using the estimates
$F(x)\ll_{k,m,r,\varepsilon}q^{\nu(r)+\varepsilon}x^{1-\frac{1}{kr}},\quad G(x)\ll_{m,r,\varepsilon}q^{\nu(r)+\varepsilon}x^{1-\frac{1}{r}},\quad|f(n)|\leq d_{k}(n)\ll_{k,\varepsilon}n^{\varepsilon},\quad|g(n)|\leq 1,$
we obtain
(2.2) $\sum_{n\leq x}(f\ast g)(n)\ll_{k,m,r,\varepsilon}q^{\nu(r)+\varepsilon}x^{1-\frac{1}{(k+1)r}}+q^{2\nu(r)+\varepsilon}x^{1-\frac{2}{(k+1)r}}.$
In view of the trivial bound
$\sum_{n\leq x}(f\ast g)(n)\ll_{k,\varepsilon}x^{1+\varepsilon},$
we may assume without loss of generality that $x\geq q^{(k+1)r\nu(r)}$. Hence the first term on the right-hand side of (2.2) majorises the second, so we get
$\sum_{n\leq x}(\chi_{1}\ast\cdots\ast\chi_{k+1})(n)\ll_{k,m,r,\varepsilon}q^{\nu(r)+\varepsilon}x^{1-\frac{1}{(k+1)r}}.$
This completes the proof. ∎
Proposition 2.4.
Let $\chi$ be a quadratic character modulo $q$, not necessarily primitive, with conductor not dividing $m$. Then there is a prime $p\equiv 1\pmod{m}$ such that $\chi(p)=-1$ and
$p\ll_{m,\varepsilon}\begin{cases}q^{\varepsilon}&\text{if GLH holds for Dirichlet $L$-functions},\\ q^{\frac{\varphi(m)}{4\sqrt{e}}+\varepsilon}&\text{otherwise}.\end{cases}$
Proof.
Note that if $a^{\flat}(n)\neq 0$ then for every prime $p\mid n$ we have either $p\equiv 1\pmod{m}$ or $2\mid v_{p}(n)$. Hence if $a^{\flat}(n)\chi(n)<0$ then there exists $p\mid n$ such that $p\equiv 1\pmod{m}$ and $\chi(p)=-1$.
Fix a small $\delta>0$, and suppose there is no such prime $\leq y=q^{\delta+\alpha e^{\delta/2}}$, where
$\alpha\coloneq\begin{cases}0&\text{if GLH holds for Dirichlet $L$-functions},\\ \frac{\varphi(m)}{4\sqrt{e}}&\text{otherwise}.\end{cases}$
Then $a^{\flat}(n)(1-\chi(n))=0$ whenever $n$ is coprime to $q$ and $y$-smooth. Considering $x\in\bigl[y^{1+\delta},y^{2}\bigr]$, we have
(2.3) $\sum_{n\leq x}a^{\flat}(n)\chi(n)$ $=\sum_{\begin{subarray}{c}n\leq x\\ (n,q)=1\end{subarray}}a^{\flat}(n)-\sum_{\begin{subarray}{c}n\leq x\\ (n,q)=1\end{subarray}}a^{\flat}(n)(1-\chi(n))$
$=\sum_{\begin{subarray}{c}n\leq x\\ (n,q)=1\end{subarray}}a^{\flat}(n)-\sum_{\begin{subarray}{c}y<p\leq x\\ p\nmid q\end{subarray}}a^{\flat}(p)\sum_{\begin{subarray}{c}n\leq x/p\\ (n,q)=1\end{subarray}}a^{\flat}(n)(1-\chi(pn))$
$\geq\sum_{\begin{subarray}{c}n\leq x\\ (n,q)=1\end{subarray}}a^{\flat}(n)-2\sum_{y<p\leq x}a(p)\sum_{\begin{subarray}{c}n\leq x/p\\ (n,q)=1\end{subarray}}a^{\flat}(n).$
By a classical result going back to Weber [23, §194], we have
$\sum_{n\leq x}a(n)=\kappa x+O\bigl(x^{\theta}\bigr)\quad\text{for all }x>0,$
where $\kappa>0$ and $\theta<1$ depend only on $m$. We may assume without loss of generality that $\theta>\frac{1}{3}$. Then applying a sieve as in the proof of Lemma 2.3, we derive
(2.4) $\sum_{n\leq x}a^{\flat}(n)=\kappa^{\flat}x+O\bigl(x^{\theta}\bigr)\quad\text{for all }x>0,$
where $\kappa^{\flat}=\kappa\prod_{\mathfrak{p}\in T}(1-N(\mathfrak{p})^{-1})>0$.
Similarly, to sieve out primes dividing $q$, let $b_{q}(n)$ be the coefficients satisfying
$\prod_{\begin{subarray}{c}\mathfrak{p}\mid q\mathcal{O}_{K}\\ \mathfrak{p}\notin T\end{subarray}}(1-N(\mathfrak{p})^{-s})=\sum_{n=1}^{\infty}\frac{b_{q}(n)}{n^{s}}.$
Then
$\sum_{\begin{subarray}{c}n\leq x\\ (n,q)=1\end{subarray}}a^{\flat}(n)$ $=\sum_{n\leq x}\sum_{d\mid n}b_{q}(d)a^{\flat}(n/d)=\sum_{d=1}^{\infty}b_{q}(d)\sum_{n\leq x/d}a^{\flat}(n)=\sum_{d=1}^{\infty}b_{q}(d)\left(\frac{\kappa^{\flat}x}{d}+O\bigl((x/d)^{\theta}\bigr)\right)$
$=\kappa^{\flat}A(q)x+O\bigl(B(q)x^{\theta}\bigr),$
where
$A(q)=\sum_{d=1}^{\infty}\frac{b_{q}(d)}{d}=\prod_{\begin{subarray}{c}\mathfrak{p}\mid q\mathcal{O}_{K}\\ \mathfrak{p}\notin T\end{subarray}}(1-N(\mathfrak{p})^{-1})\geq\left(\frac{\varphi(q)}{q}\right)^{\varphi(m)}\gg_{m}(\log\log{q})^{-\varphi(m)}$
and
$B(q)=\sum_{d=1}^{\infty}\frac{|b_{q}(d)|}{d^{\theta}}=\prod_{\begin{subarray}{c}\mathfrak{p}\mid q\mathcal{O}_{K}\\ \mathfrak{p}\notin T\end{subarray}}\bigl(1+N(\mathfrak{p})^{-\theta}\bigr)\ll_{m,\varepsilon}q^{\varepsilon}.$
Returning to (2.3), we split the $p$ sum over the ranges $y<p\leq z$ and $z<p\leq x$, where $z=(\frac{A(q)}{B(q)})^{\frac{1}{1-\theta}}x\gg_{\varepsilon}x^{1-\varepsilon}$. (Note that $z>y$ when $q$ is sufficiently large.) For the small primes, we have
$2\sum_{y<p\leq z}a(p)\sum_{\begin{subarray}{c}n\leq x/p\\ (n,q)=1\end{subarray}}a^{\flat}(n)$ $=2\sum_{y<p\leq z}a(p)\left(\frac{\kappa^{\flat}A(q)x}{p}+O\bigl(B(q)(x/p)^{\theta}\bigr)\right)$
$=2\kappa^{\flat}A(q)x\sum_{y<p\leq z}\frac{a(p)}{p}+O\!\left(B(q)x^{\theta}\sum_{y<p\leq z}|a(p)|p^{-\theta}\right).$
Recall that $a(p)=\sum_{\xi}\xi(p)$, where $\xi$ ranges over the primitive characters of conductor dividing $m$. By the prime number theorem in arithmetic progressions, we have
$\sum_{y<p\leq z}\frac{a(p)}{p}=\log\!\left(\frac{\log{z}}{\log{y}}\right)+O_{m}\bigl((\log{y})^{-1}\bigr),$
and using that $|a(p)|\leq\varphi(m)$, we have
$\frac{B(q)}{A(q)}\sum_{y<p\leq z}|a(p)|(x/p)^{\theta}\ll_{m}\frac{B(q)}{A(q)}x^{\theta}\frac{z^{1-\theta}}{\log{x}}=\frac{x}{\log{x}}.$
Thus the small prime sum is
$2\kappa^{\flat}A(q)x\left(\log\!\left(\frac{\log{z}}{\log{y}}\right)+O_{m}\bigl((\log{x})^{-1}\bigr)\right).$
For the large primes we change order of summation:
$2\sum_{z<p\leq x}a(p)$ $\sum_{\begin{subarray}{c}n\leq x/p\\ (n,q)=1\end{subarray}}a^{\flat}(n)\leq 2\sum_{\begin{subarray}{c}n\leq x/z\\ (n,q)=1\end{subarray}}a^{\flat}(n)\sum_{p\leq x/n}a(p).$
Applying the prime number theorem in arithmetic progressions, the inner sum is
$\sum_{p\leq x/n}a(p)=\frac{x}{n\log{x}}\left(1+O_{m}\!\left(\frac{1+\log(x/z)}{\log{x}}\right)\right),$
so we get
$\frac{2x}{\log{x}}\left(1+O\!\left(\frac{1+\log(x/z)}{\log{x}}\right)\right)\sum_{\begin{subarray}{c}n\leq x/z\\ (n,q)=1\end{subarray}}\frac{a^{\flat}(n)}{n}.$
Now, applying partial summation to (2.4), we have
$\sum_{n\leq x}\frac{a^{\flat}(n)}{n}=\kappa^{\flat}\log{x}+c+O\bigl(x^{\theta-1}\bigr)\quad\text{for all }x>0,$
for a certain constant $c$ depending only on $m$, and hence
$\sum_{\begin{subarray}{c}n\leq x\\ (n,q)=1\end{subarray}}\frac{a^{\flat}(n)}{n}$ $=\sum_{d=1}^{\infty}\frac{b_{q}(d)}{d}\sum_{n\leq x/d}\frac{a^{\flat}(n)}{n}=\sum_{d=1}^{\infty}\frac{b_{q}(d)}{d}\bigl(\kappa^{\flat}\log(x/d)+c+O\bigl((x/d)^{\theta-1}\bigr)\bigr)$
$=A(q)(\kappa^{\flat}\log{x}+c)-\kappa^{\flat}\sum_{d=1}^{\infty}\frac{b_{q}(d)\log{d}}{d}+O\bigl(B(q)x^{\theta-1}\bigr)$
$=A(q)\left(\kappa^{\flat}\log{x}+c+\kappa^{\flat}\sum_{\begin{subarray}{c}\mathfrak{p}\mid q\mathcal{O}_{K}\\ \mathfrak{p}\notin T\end{subarray}}\frac{\log{N(\mathfrak{p})}}{N(\mathfrak{p})-1}\right)+O\bigl(B(q)x^{\theta-1}\bigr)$
$=\kappa^{\flat}A(q)\bigl(\log{x}+O(\log\log{q})\bigr)+O\bigl(B(q)x^{\theta-1}\bigr).$
Replacing $x$ by $x/z=(B(q)/A(q))^{\frac{1}{1-\theta}}$, we see that the large prime sum is at most
$2\kappa^{\flat}A(q)\frac{x}{\log{x}}\left(1+O\!\left(\frac{1+\log(x/z)}{\log{x}}\right)\right)(\log(x/z)+O(\log\log{q}))$
$=2\kappa^{\flat}A(q)x\left(\frac{\log(x/z)}{\log{x}}+O\!\left(\frac{\log\log{q}}{\log{q}}\right)\right).$
Combining with the small prime sum, we get
$\leq 2\kappa^{\flat}A(q)x\left(\log\!\left(\frac{\log{x}}{\log{y}}\right)+\log\!\left(\frac{\log{z}}{\log{x}}\right)+\frac{\log(x/z)}{\log{x}}+O\!\left(\frac{\log\log{q}}{\log{q}}\right)\right)$
$\leq 2\kappa^{\flat}A(q)x\left(\log\!\left(\frac{\log{x}}{\log{y}}\right)+O\!\left(\frac{\log\log{q}}{\log{q}}\right)\right),$
and so altogether we have
$\sum_{n\leq x}a^{\flat}(n)\chi(n)\geq\kappa^{\flat}A(q)x\left(1-2\log\!\left(\frac{\log{x}}{\log{y}}\right)+O\!\left(\frac{\log\log{q}}{\log{q}}\right)\right).$
Now set $x=y^{e^{\frac{1}{2}(1-\delta)}}$ . Then
$1-2\log\!\left(\frac{\log{x}}{\log{y}}\right)+O\!\left(\frac{\log\log{q}}{\log{q}}\right)=\delta+O\!\left(\frac{\log\log{q}}{\log{q}}\right).$
This is at least $\delta/2$ for sufficiently large $q$, so applying Lemma 2.3 with $r=2$ when GLH holds and $r>\frac{\varphi(m)}{4\delta\exp((1-\delta)/2)}$ otherwise, we have
$(\log\log{q})^{-\varphi(m)}$ $\ll A(q)\ll_{\delta}\frac{1}{x}\left|\sum_{n\leq x}a^{\flat}(n)\chi(n)\right|\ll_{r,\varepsilon}q^{\nu(r)+\varepsilon}x^{-\frac{1}{\varphi(m)r}}$
$=q^{\varepsilon-\frac{\delta\exp((1-\delta)/2)}{\varphi(m)r}}\begin{cases}1&\text{if GLH holds},\\ q^{\frac{1}{4r^{2}}}&\text{otherwise}.\end{cases}$
Choosing $\varepsilon$ sufficiently small, we obtain a contradiction for sufficiently large $q$.
Hence for $q\gg_{m,\delta}1$ there exists a prime $p\leq y=q^{\delta+\alpha e^{\delta/2}}$ satisfying $p\equiv 1\pmod{m}$ and $\chi(p)=-1$. By Dirichlet’s theorem, there is always a suitable $p\ll_{q}1$, so we can choose an implied constant to cover the small values of $q$. Since $\delta$ was arbitrary, we thus have $p\ll_{m,\varepsilon}q^{\alpha+\varepsilon}$. ∎
Theorem 2.5.
For $i=1,\ldots,r$, let $\chi_{i}\pmod{q_{i}}$ be a quadratic character, and let $\epsilon_{i}\in\{\pm 1\}$. Assume that no non-empty product of $\chi_{i}$s has conductor dividing $m$. Then there is a squarefree positive integer $n$ with at most $r$ prime factors $p_{j}$, each satisfying
$p_{j}\equiv 1\!\!\!\!\pmod{m}\quad\text{and}\quad p_{j}\ll_{m,\varepsilon}\begin{cases}(q_{1}\cdots q_{r})^{\varepsilon}&\text{if GLH holds for Dirichlet $L$-functions},\\ (q_{1}\cdots q_{r})^{\frac{\varphi(m)}{4\sqrt{e}}+\varepsilon}&\text{otherwise},\end{cases}$
such that $\chi_{i}(n)=\epsilon_{i}$ for all $i=1,\ldots,r$.
Proof.
This generalises [4, Lemma 4], and the proof is identical, substituting Proposition 2.4 in place of [4, Lemma 3]. ∎
3. Conditional analytic estimates
Next we derive an analogue of Theorem 2.5 for higher-order characters that will be useful in the proof of Theorem 1.4.
Lemma 3.1.
[…]
Proof.
[…]
Hence it suffices to find $P\nmid q$ with $P\equiv 1\pmod{\ell}$ such that $P$ does not split completely in $L$. Equivalently, we seek an unramified degree $1$ prime $\mathfrak{p}$ of $K$ such that the Artin symbol $\left(\frac{L/K}{\mathfrak{p}}\right)$ is non-trivial. By [1, Theorem 3.1(2)], there is such a $\mathfrak{p}$ satisfying
$N_{K/\mathbb{Q}}\mathfrak{p}\leq(1+o(1))\log^{2}\Delta_{L},$
where $\Delta_{L}$ is the absolute discriminant of $L/\mathbb{Q}$, and $o(1)$ is a quantity bounded by an absolute constant and tending to 0 as $\Delta_{L}\to\infty$.
[…]
$\Delta_{L}=\Delta_{K}^{[L:K]}N_{K/\mathbb{Q}}(\operatorname{Disc}(L/K))=\ell^{\ell(\ell-2)}\prod_{p\mid q}p^{(\ell-1)^{2}},$
and thus we can take
$P=N_{K/\mathbb{Q}}\mathfrak{p}\ll\log^{2}\Delta_{L}\ll\ell^{4}\log^{2}\left(\ell\prod_{p\mid q}p\right).$
Theorem 3.2.
[…]
  • $\chi$ has order dividing $\ell$;
  • every prime $p\mid d$ satisfies $p\ll\ell^{4}\log^{2}(\ell q_{1}\cdots q_{r})$;
  • $\chi(q_{j})=\epsilon_{j}$ for $j=1,\ldots,r$.
Proof.
[…]
Therefore the $w_{i}$ span $\mathbb{F}_{\ell}^{r}$, so there are indices $i_{1},\ldots,i_{r}\leq n$ such that $\{w_{i_{1}},\ldots,w_{i_{r}}\}$ is a basis. Let $b=(b_{1},\ldots,b_{r})\in\mathbb{F}_{\ell}^{r}$ be the vector such that $\epsilon_{j}=\zeta^{b_{j}}$, and write $b=a_{1}w_{i_{1}}+\cdots+a_{r}w_{i_{r}}$. Then the character
$\chi=\prod_{\begin{subarray}{c}1\leq k\leq r\\ a_{k}\neq 0\end{subarray}}\chi_{i_{k}}^{a_{k}}$
fulfills the requirements of the lemma. ∎
4. Results on Diophantine equations
Lemma 4.1.
Let $f\in\mathbb{Z}[x]$ be a squarefree polynomial of degree at least $3$. Then $f(n)$ is a square for at most finitely many $n\in\mathbb{Z}$.
Proof.
This follows from Siegel’s theorem on integral points [18]. ∎
We also provide a simple proof of this result for polynomials of even degree with square leading coefficient:
Lemma 4.2.
Suppose $f\in\mathbb{Z}[x]$ has even degree and leading coefficient $a^{2}$ for some $a\in\mathbb{Z}$. Then either $f=g^{2}$ for some $g\in\mathbb{Z}[x]$ or $f(n)$ is a square for at most finitely many $n\in\mathbb{Z}$.
Proof.
Let $d=\deg{f}$. The conclusion is clear if $f$ is constant, so we may assume that $d>0$. We claim that there are polynomials $s,r\in\mathbb{Z}[x]$ with $\deg{r}<\frac{d}{2}$ such that
$(2a)^{2d-2}f(x)=s(x)^{2}+r(x).$
To see this, write $f(x)=\sum_{i=0}^{d}f_{i}x^{d-i}$ and $s(x)=\sum_{i=0}^{\frac{d}{2}}s_{i}x^{\frac{d}{2}-i}$, where we define $s_{0}=2^{d-1}a^{d}$, and recursively for $i=1,2,\ldots,\frac{d}{2}$,
$s_{i}=(2a)^{d-2}f_{i}-(2a)^{-d}\sum_{j=1}^{i-1}s_{j}s_{i-j}.$
Then it is clear that $r(x)\coloneq(2a)^{2d-2}f(x)-s(x)^{2}$ has degree less than $\frac{d}{2}$, and by induction we see that $s_{i}\in(2a)^{d-2i}\mathbb{Z}$ for every $i>0$, so $s,r\in\mathbb{Z}[x]$.
Since $s$ has degree $\frac{d}{2}$, for large $x$ we have |r(x)1|<2|s(x)||r(x)-1|<2|s(x)|, and it follows that $(2a)^{2d-2}f(x)$ is strictly between $(|s(x)|-1)^{2}$ and $(|s(x)|+1)^{2}$.
If $r$ is not identically 0 then it has at most $\frac{d}{2}-1$ zeros, and therefore $f(x)$ has at most finitely many square values. Otherwise, if $r$ is identically 0 then Gauss’ Lemma implies that $f=g^{2}$ for some $g\in\mathbb{Z}[x]$. ∎
Lemma 4.3.
Let $N\in\mathbb{N}$ with $\sqrt{N}\notin\mathbb{N}$. Then there are infinitely many primes $p$ such that the Kronecker symbol $\left(\frac{p}{N}\right)$ equals 1-1.
Proof.
Since $N$ is not a square, $\left(\frac{\cdot}{4N}\right)$ is a non-principal quadratic character modulo $4N$, and hence we have $\sum_{n=1}^{4N}\left(\frac{n}{4N}\right)=0$. Since $\left(\frac{1}{4N}\right)=1$, it follows that there is some $n\leq 4N$ with $\left(\frac{n}{4N}\right)=-1$.
We conclude with a version of Euclid’s argument. Let $p_{1},\ldots,p_{k}$ be a (possibly empty) sequence of distinct primes satisfying $\left(\frac{p_{i}}{4N}\right)=-1$ and $p_{i}\nmid n$, and consider the number $t=4Np_{1}\cdots p_{k}+n$. Since $t\equiv n\pmod{4N}$, we have $\left(\frac{t}{4N}\right)=-1$, and by multiplicativity it follows that there is a prime $p\mid t$ with $\left(\frac{p}{4N}\right)=-1$. By hypothesis $(t,n)=(4Np_{1}\cdots p_{k},n)=1$, so $p\nmid n$, and it follows that $p\notin\{p_{1},\ldots,p_{k}\}$. ∎
Lemma 4.4.
[…]
Proof.
[…]
$\frac{1}{[L:K]}\left(1-\frac{1}{[L(\zeta_{16}):L]}\right)\prod_{i=1}^{n}\left(1-\frac{1}{[L(\sqrt[\ell]{\alpha_{i}}):L]}\right)\prod_{p\mid m}\left(1-\frac{1}{[L(\sqrt[p]{\epsilon}):L]}\right),$
and this is positive provided that $\zeta_{16}$, $\sqrt[\ell]{\alpha_{i}}$, and $\sqrt[p]{\epsilon}$ for $p\mid m$ are not contained in $L$.
To that end, first note that none of $\sqrt[4]{\epsilon}$, $\sqrt[\ell]{\alpha_{i}}$, or $\sqrt[p]{\epsilon}$ for $p\mid m$ is contained in $K(\zeta_{r})$ for any $r$, since they have both real and non-real conjugates, while $K(\zeta_{r})$ is an abelian extension of $\mathbb{Q}$. Combining this with a consideration of degrees, we find that $\sqrt[\ell]{\alpha_{i}}$ and $\sqrt[p]{\epsilon}$ are not contained in $L$.
Next we aim to show that $\zeta_{16}\notin L$. Suppose otherwise, and let $d\mid m$ be the smallest number such that $\zeta_{16}\in K(\zeta_{8d},\sqrt[8]{\epsilon})$. Then for any prime $p\mid d$, $\zeta_{16}$ is not contained in $M\coloneq K(\zeta_{8d/p},\sqrt[8]{\epsilon})$ but is contained in $M(\zeta_{p^{v_{p}(d)}})$. Therefore $\operatorname{Disc}(M(\zeta_{16})/M)\mid\operatorname{Disc}(M(\zeta_{p^{v_{p}(d)}})/M)$, and so $p=2$. Since $m$ is odd this cannot happen, so we must have $d=1$ and $\zeta_{16}\in K(\zeta_{8},\sqrt[8]{\epsilon})$.
Suppose $K\neq\mathbb{Q}(\sqrt{2})$, and let $F=K(\zeta_{8})$. Let $k\in\{2,4\}$ be the smallest power of $2$ such that $\sqrt[k]{\epsilon}\notin F$. Since $K\neq\mathbb{Q}(\sqrt{2})$, we have $\operatorname{Disc}{K}\nmid 16$, so $K\not\subset\mathbb{Q}(\zeta_{16})$, and hence $\zeta_{16}\notin F$. Therefore $F(\zeta_{16})=F(\sqrt{\zeta_{8}})$ is a quadratic extension of $F$ contained in $F(\sqrt[8]{\epsilon})$. Kummer theory implies that $F(\sqrt[k]{\epsilon})$ is the unique such extension, and it follows that $\sqrt[k]{\epsilon}\zeta_{16}\in F$. In particular, $\sqrt{\epsilon}\in K(\zeta_{16})$. Since $K\cap\mathbb{Q}(\zeta_{16})=K\cap\mathbb{Q}(\sqrt{2})=\mathbb{Q}$, an exercise in Galois theory shows that
$\operatorname{Gal}(K(\zeta_{16})/K)\cong\operatorname{Gal}(\mathbb{Q}(\zeta_{16})/\mathbb{Q})\cong(\mathbb{Z}/16\mathbb{Z})^{\times}\cong C_{4}\times C_{2},$
so $K(\zeta_{16})$ contains three quadratic extensions of $K$: $K(\sqrt{2})$, $K(i)$, and $K(i\sqrt{2})$. Among these, only $K(\sqrt{2})$ has a real embedding, so we must have $K(\sqrt{\epsilon})=K(\sqrt{2})\subset F$, meaning $k=4$. Supposing that $\sqrt[4]{\epsilon}\zeta_{16}=a+bi\in F=K(\sqrt{2},i)$ and taking the norm to $K(\sqrt{2})$, we obtain $\sqrt{\epsilon}=a^{2}+b^{2}$. This contradicts the fact that $\sqrt{\epsilon}$ is not totally positive, so $\zeta_{16}\notin L$.
It remains only to consider $K=\mathbb{Q}(\sqrt{2})$. In this case the only quadratic extension of $K$ contained in $K(\zeta_{16})=K\bigl(\sqrt{2+\sqrt{2}},i\bigr)$ with a real embedding is $K\bigl(\sqrt{2+\sqrt{2}}\bigr)$, which cannot equal $K(\sqrt{\epsilon})$ since $N(\epsilon)=-1$. It follows that $K(\zeta_{16},\sqrt[8]{\epsilon}):K(\zeta_{8})]=16$, so again $\zeta_{16}\notin L$.
Lemma 4.5.
  • (i)
    There are no solutions to $x^{2}+1=y^{r}$ in positive integers with $r>1$.
  • (ii)
    If $x,y,r$ are positive integers satisfying $x^{2}+1=2y^{r}$ and $r>2$ then either $x=y=1$ or $(x,y,r)=(239,13,4)$.
Proof.
These are results of Lebesgue [9] and Störmer [20, §3], respectively. ∎
5. Proof of main results
Proof of Theorem 1.4.
Consider $m\in\{1,2\}$ and $f\in\mathrm{GEM}(1,m)$. By Gauss’ Lemma, we can write $f=c\prod_{i}f_{i}^{n_{i}}$, where $c\in\mathbb{Z}$ and $f_{i}\in\mathbb{Z}[x]$ is irreducible with content $1$. It is easy to see from Definition 1.1 that $F(x)=c^{2}\prod_{i}(f_{i}(0)f_{i}(x))$ is a $\mathrm{GEM}(1,m)$ polynomial satisfying $\mathrm{GEM}_{2}(F;1,m)=\mathrm{GEM}_{2}(f;1,m)$. Thus, replacing $f$ by $F$ if necessary, we may assume without loss of generality that $f$ is squarefree and $f(0)$ is a square.
Write $\mathrm{GEM}_{2}(f;1,m)=(p_{n})_{n\geq 1}$, and suppose that it contains all primes except for a finite set $\{q_{1},\ldots,q_{r}\}$. (Note that the exceptional set must contain all primes dividing $mf(0)$, but could otherwise be empty.) Let $x>0$ be a large real number, let $p=p_{n+1}$ be the prime $\leq x$ that occurs last in the sequence, and set $N=|f(p_{1}\cdots p_{n})|$. Then
$N=q_{1}^{e_{1}}\cdots q_{r}^{e_{r}}p^{e},$
for some exponents $e_{i}\geq 0$ and $e>0$.
Suppose first that $\deg{f}\geq 3$. Then applying Lemma 4.1 to both $f$ and f-f, we see that $N$ is not a square when $x$ is sufficiently large, so by Lemma 4.3 there is a prime $P$ exceeding $q_{1},\ldots,q_{r}$ such that $\left(\frac{P}{N}\right)=-1$. Next we apply Theorem 2.5 to the characters
$\left(\frac{\cdot}{q_{1}}\right),\ldots,\left(\frac{\cdot}{q_{r}}\right),\left(\frac{\cdot}{p}\right),\left(\frac{-4}{\cdot}\right).$
This produces a squarefree positive integer $d\equiv 1\pmod{4}$ such that
$\left(\frac{d}{q_{i}}\right)=\left(\frac{P}{q_{i}}\right),\quad\left(\frac{d}{p}\right)=\left(\frac{P}{p}\right)$
and every prime factor of $d$ is $O_{\varepsilon}\bigl((q_{1}\cdots q_{r}p)^{\frac{1}{4\sqrt{e}}+\varepsilon}\bigr)$. Since $\frac{1}{4\sqrt{e}}<1$, this is less than $p$ for sufficiently large $x$, whence $d\mid p_{1}\cdots p_{n}$.
By multiplicativity of the Kronecker symbol, it follows that $\left(\frac{d}{N}\right)=\left(\frac{P}{N}\right)=-1$. However, by our construction of $d$, we have $N\equiv\pm f(0)\pmod{d}$, so by quadratic reciprocity,
$\left(\frac{d}{N}\right)=\left(\frac{N}{d}\right)=\left(\frac{\pm f(0)}{d}\right)=1.$
This is a contradiction, completing the proof when $\deg{f}\geq 3$.
Next suppose $\deg{f}=2$, and write $f=ax^{2}+bx+c$ with $c$ a square, and $D\coloneq b^{2}-4ac\neq 0$. If $a$ is also a square then it follows from Lemma 4.2 that $|f|$ has at most finitely many square values, so the above proof goes through. Hence we may assume that $\sqrt{a}\notin\mathbb{Q}$. If $a<0$ then, choosing $p_{i}\equiv 3\pmod{4}$, for large enough $x$ it follows that
$\left(\frac{N}{p_{i}}\right)=\left(\frac{-f(p_{1}\cdots p_{n})}{p_{i}}\right)=\left(\frac{-c}{p_{i}}\right)=-1.$
Thus $N$ is not a square, and again the proof goes through. Hence we may assume that $a>0$.
Suppose $f(u)=v^{2}$ for some $u,v\in\mathbb{Z}$. Completing the square, we have
(5.1) $(2au+b)^{2}-4av^{2}=D.$
Let $K=\mathbb{Q}(\sqrt{a})$, with ring of integers $\mathcal{O}_{K}$ and fundamental unit $\epsilon\in\mathcal{O}_{K}^{\times}$, and let $\sigma:K\to K$ denote the non-trivial automorphism. By the theory of Pell’s equation, there are elements $\rho_{1},\ldots,\rho_{s}\in\mathcal{O}_{K}$ such that
  • $N(\rho_{i})=D$ for each $i$;
  • $\rho_{i}/\rho_{j}\notin\mathcal{O}_{K}^{\times}$ for $i\neq j$;
  • for any solution to (5.1), there exist $i\in\{1,\ldots,s\}$ and $k\in\mathbb{Z}$ such that $N(\epsilon)^{k}=1$ and
    $2au+b+v\sqrt{4a}=\pm\rho_{i}\epsilon^{k}.$
Returning to the setup above, we wish to rule out solutions to (5.1) with $u=p_{1}\cdots p_{n}$. If the desired conclusion is false then we can generate infinitely many such solutions by increasing $x$. In each solution we may assume that $k\geq 0$ by applying $\sigma$ if necessary, and by the pigeonhole principle we can choose infinitely many solutions with a fixed value of $\rho_{i}=\rho$ and a fixed $\pm$ sign. Replacing $\rho$ by $\rho$ if necessary, we thus obtain
$4ap_{1}\cdots p_{n}=\rho\epsilon^{k}+\rho^{\sigma}\epsilon^{-k}-2b$
for infinitely many $n$ (with $k$ depending on $n$, but $a,b,\rho,\epsilon$ fixed). Set $\rho_{0}=b+\sqrt{4ac}\in\mathcal{O}_{K}$, and note that $N(\rho_{0})=D$ and $\rho_{0}+\rho_{0}^{\sigma}=2b$, so the above can be written as
$4ap_{1}\cdots p_{n}=\rho\epsilon^{k}-\rho_{0}+\rho^{\sigma}\epsilon^{-k}-\rho_{0}^{\sigma}=\frac{(\rho\epsilon^{k}-\rho_{0})(\rho\epsilon^{k}-\rho_{0}^{\sigma})}{\rho\epsilon^{k}}.$
Thus if $p_{i}$ is a prime not dividing $D$ then for each prime $\mathfrak{p}\mid p_{i}\mathcal{O}_{K}$ we have
$\epsilon^{k}\equiv\rho^{-1}\rho_{0}\text{ or }\rho^{-1}\rho_{0}^{\sigma}\pmod{\mathfrak{p}}$.
[…]
Hence we have $\rho\in\rho_{0}\mathcal{O}_{K}^{\times}\cup\rho_{0}^{\sigma}\mathcal{O}_{K}^{\times}$, and replacing $\rho_{0}$ by $\rho_{0}^{\sigma}$ if necessary we may assume that $\rho\in\rho_{0}\mathcal{O}_{K}^{\times}$. Further, replacing $k$ by a suitable translate, we may assume that $\rho=\pm\rho_{0}$, so we obtain
$4ap_{1}\cdots p_{n}=\epsilon^{-k}(\epsilon^{k}\mp 1)(\pm\rho_{0}\epsilon^{k}-\rho_{0}^{\sigma}).$
[…]
$4ap_{1}\cdots p_{n}=\epsilon^{-k}(\epsilon^{k}-1)(\rho_{0}\epsilon^{k}-\rho_{0}^{\sigma}).$
[…]
Suppose now that $\rho_{0}^{-1}\rho_{0}^{\sigma}\in\mathcal{O}_{K}^{\times}$. Then it follows that $|D|$ is a square and $\rho_{0}\in\sqrt{|D|}\mathcal{O}_{K}^{\times}$, so we have $|\rho_{0}|=\sqrt{|D|}\epsilon^{j}$ for some $j\in\mathbb{Z}$. Thus, up to a constant factor, we need to analyse the reduction mod $\mathfrak{p}$ of $(\epsilon^{k}\pm 1)(\epsilon^{k+2j}\pm\operatorname{sgn}{D})$. Applying Lemma 4.4 again, we see that the $(\epsilon^{k}+1)(\epsilon^{k+2j}+1)$ case cannot occur, and in the remaining cases we conclude that $m$ must divide one of $k$ or $k+2j$. Since $m$ is odd, this still suffices to imply that $\mathfrak{q}^{2}\mid p_{1}\cdots p_{n}\mathcal{O}_{K}$, completing the proof when $\deg{f}=2$.
Finally suppose $\deg{f}=1$, and write $f=ax+b$ with $b$ a square. The case when $a<0$ can be handled by reducing modulo a prime $p_{i}\equiv 3\pmod{4}$ as in the proof for degree $2$ above, so we may assume that $a>0$. Let $g=(a,b)$, and set $a^{\prime}=a/g$, $b^{\prime}=b/g$. By our construction, $p_{1}\cdots p_{n}$ is coprime to $b$, and it follows that $N^{\prime}\coloneq\frac{N}{g}=a^{\prime}p_{1}\cdots p_{n}+b^{\prime}$ is coprime to $b^{\prime}$. Suppose $b^{\prime}N^{\prime}$ is not an $8$th power, and note that this must be the case if $2\mid a^{\prime}$ and $32\nmid a^{\prime}$. Let $m\in\{2,4,8\}$ be the least power of $2$ such that $b^{\prime}N^{\prime}$ is not an $m$th power, and set $y=(b^{\prime}N^{\prime})^{\frac{2}{m}}\in\mathbb{N}$. Then $y$ is not a square, so by Lemma 4.3 there is a prime $P$ exceeding $q_{1},\ldots,q_{r}$ for which $\left(\frac{P}{y}\right)=-1$. Applying Theorem 2.5 as before, we obtain a squarefree number $d$ such that
$\left(\frac{d}{q_{i}}\right)=\left(\frac{P}{q_{i}}\right),\quad\left(\frac{d}{p}\right)=\left(\frac{P}{p}\right),\quad\left(\frac{-4}{d}\right)=1\text{ when }m=2,$
and every prime factor of $d$ is $O_{m,\varepsilon}\bigl((q_{1}\cdots q_{r}p)^{\frac{\varphi(m)}{4\sqrt{e}}+\varepsilon}\bigr)$ and congruent to $1\pmod{m}$. Since $\frac{\varphi(m)}{4\sqrt{e}}<1$, it follows that $d\mid p_{1}\cdots p_{n}$ when $x$ is sufficiently large. For $p_{i}\equiv 1\pmod{m}$, Euler’s criterion implies
$\left(\frac{y}{p_{i}}\right)\equiv y^{\frac{p_{i}-1}{2}}=(b^{\prime}N^{\prime})^{\frac{p_{i}-1}{m}}\equiv\bigl((b^{\prime})^{2}\bigr)^{\frac{p_{i}-1}{m}}=\bigl((b^{\prime})^{\frac{2}{m}}\bigr)^{p_{i}-1}\equiv 1\pmod{p_{i}}$.
Since all prime factors of $b^{\prime}N^{\prime}$ are contained in $\{q_{1},\ldots,q_{r},p\}$, we thus have
$1=\left(\frac{P}{y}\right)=\left(\frac{d}{y}\right)=\left(\frac{y}{d}\right)=1,$
which is again a contradiction.
[…]
Assume that $x$ is large enough to ensure that $\ell\mid p_{1}\cdots p_{n}$. Then we have $N\equiv b\pmod{\ell}$ but $N\not\equiv b\pmod{\ell^{2}}$ since $\ell^{2}\nmid ap_{1}\cdots p_{n}$. Therefore, if $\xi\pmod{\ell^{2}}$ is a character of order $\ell$ then $\xi(N)\neq 1$. Assuming ERH, we can apply Theorem 3.2 to choose a character $\chi$ of squarefree conductor $d$ and order dividing $\ell$ such that
$\chi(p)=\xi(p)\quad\text{and}\quad\chi(q_{i})=\xi(q_{i})\text{ for }i=1,\ldots,r,$
and every prime factor of $d$ is $O(\ell^{4}\log^{2}(\ell q_{1}\cdots q_{r}p))$. For sufficiently large $x$ this is less than $p$, and therefore $d\mid p_{1}\cdots p_{n}$, so that $\chi(N)=\chi(b)=1$. However, by construction we have
$\chi(N)=\chi(q_{1})^{e_{1}}\cdots\chi(q_{r})^{e_{r}}\chi(p)^{e}=\xi(q_{1})^{e_{1}}\cdots\xi(q_{r})^{e_{r}}\xi(p)^{e}=\xi(N)\neq 1.$
This is a contradiction, and that completes the proof of Theorem 1.4. ∎
Proof of Theorems 1.6 and 1.7.
Let $m\in\mathbb{N}$ and $c\in S(m)$. We may assume that $m\neq 2$, and that $m=1\implies c\in\{-5\}\cup\{1-2^{k}:k\in\mathbb{N}\}$, since Theorem 1.6 and all other cases of Theorem 1.7 for $m\leq 2$ follow from Theorem 1.4. As in the proof above we suppose that $\mathrm{GEM}_{2}(\Phi_{m}(cx);1,m)=(p_{n})_{n\geq 1}$ contains all primes $\equiv 1\pmod{m}$ except for a finite (possibly empty) set $\{q_{1},\ldots,q_{r}\}$. Let $\mu$ be the largest prime factor of $m$ if $m>1$, and set $\mu=1$ if $m=1$. Let $x>0$ be a large real number, let $p=p_{n+1}$ be the prime $\leq x$ that occurs last in the sequence, and set $M=|\Phi_{m}(cp_{1}\cdots p_{n})|$. Then we have
$M=\mu^{\delta}q_{1}^{e_{1}}\cdots q_{r}^{e_{r}}p^{e}$
for some exponents $\delta\in\{0,1\}$, $e_{i}\geq 0$, and $e>0$, with $\delta=0$ when $m=1$.
Set $N=M/\mu^{\delta}$. Then we claim that $N$ is not a square for $x$ sufficiently large:
  • For $m>2$ and $\delta=0$, the claim follows from Lemma 4.2 applied to $\Phi_{m}$.
  • For $m\notin\{3,4,6\}$ and $\delta=1$, the claim follows from Lemma 4.1 applied to $\mu\Phi_{m}$.
  • For $m\in\{3,4,6\}$ and $\delta=1$, for large enough $x$ we can choose $p_{i}$ for $i\leq n$ such that $\left(\frac{\mu}{p_{i}}\right)=-1$ (viz. $p_{i}\equiv 5\pmod{8}$ for $m=4$, $p_{i}\equiv 7\pmod{12}$ for $m\in\{3,6\}$). Since $M\equiv 1\pmod{p_{i}}$, it follows that $\left(\frac{N}{p_{i}}\right)=-1$.
  • For $m=1$ and $c\in\{-5\}\cup\bigl\{1-2^{k}:k\in\mathbb{N}\bigr\}$, $2$ occurs in the sequence as either the first or second term, after which point we have $N\equiv 3\pmod{4}$.
Applying Lemma 4.3, there exists a prime $P$ exceeding $q_{1},\ldots,q_{r}$ such that $\left(\frac{P}{N}\right)=-1$. Next we apply Theorem 2.5 to the characters
$\left(\frac{\cdot}{q_{1}}\right),\ldots,\left(\frac{\cdot}{q_{r}}\right),\left(\frac{\cdot}{p}\right),$
and also $\left(\frac{-4}{\cdot}\right)$ when $4\nmid m$ and $\left(\frac{8}{\cdot}\right)$ when $m=4$. (Note that we allow $q_{i}=2$ when $m=1$, in which case $\left(\frac{\cdot}{q_{i}}\right)=\left(\frac{8}{\cdot}\right)$ is a character of conductor $8$.) This produces a squarefree $d\equiv 1\pmod{4}$ such that
$\left(\frac{d}{q_{i}}\right)=\left(\frac{P}{q_{i}}\right),\quad\left(\frac{d}{p}\right)=\left(\frac{P}{p}\right),\quad\left(\frac{d}{2}\right)=1\text{ when }m=4,$
and every prime factor of $d$ is $O_{m,\varepsilon}\bigl((q_{1}\cdots q_{r}p)^{\frac{\varphi(m)}{4\sqrt{e}}+\varepsilon}\bigr)$ and congruent to $1\pmod{m}$. When $m\in\{1,\ldots,10,12,14,18\}$ we have $\frac{\varphi(m)}{4\sqrt{e}}<1$, so for sufficiently large $x$ every prime factor of $d$ is less than $p$, whence $d\mid p_{1}\cdots p_{n}$. For general $m$ the same conclusion holds under the assumption of GLH.
By construction we have $M\equiv\pm 1\pmod{d}$ and $d\equiv 1\pmod{m}$, and thus
$1=\left(\frac{P}{N}\right)=\left(\frac{d}{N}\right)=\left(\frac{N}{d}\right)=\left(\frac{\pm\mu^{\delta}}{d}\right)=\left(\frac{d}{\mu^{\delta}}\right)=1.$
This is a contradiction, and that completes the proof. ∎
Proof of Theorem 1.8.
Let $m=2^{k}$ for some $k>1$, and write $\mathrm{GEM}_{2}(\Phi_{m}(cx);1,m)=(p_{n})_{n\geq 1}$. Let $p$ be the smallest prime congruent to $1\pmod{m}$, and suppose that $p=p_{n+1}$ for some $n\geq 0$.
Suppose $c$ is even. Then we have
$(cp_{1}\cdots p_{n})^{2^{k-1}}+1=p^{r}.$
By Lemma 4.5(i), this can only happen if $r=1$, which forces $n=0$.
On the other hand, if $c$ is odd then we have
$(cp_{1}\cdots p_{n})^{2^{k-1}}+1=2p^{r}.$
Noting that $239$ is a prime congruent to $3\pmod{4}$, Lemma 4.5(ii) implies that if $n>0$ then we must have $r\leq 2$, which is a contradiction. Thus we again have $n=0$, and either $(k,c,p,r)=(2,\pm 239,13,4)$ or $r\leq 2$. However, if $k=2$ then $p=5$, so the former case does not apply.
Hence, in all cases we have
(5.2) $c^{2^{k-1}}+1\in\bigl\{p,2p,2p^{2}\bigr\},$
and it follows that $2p^{2}>3^{2^{k-1}}$. Set $x=3^{2^{k-2}}/\sqrt{2}$. Then by [2, Theorem 1.2],
$(80\log{3}-1)2^{k-1}>80\log{2}\implies\theta(x;2^{k},1)>\frac{x}{2^{k-1}}-\frac{x}{160\log{x}}>0,$
provided that $2^{k}>80$ and $x>\exp(0.03\cdot 2^{k/2}\log^{3}(2^{k}))$, which holds for $k\geq 12$. Noting that $17\equiv 1\pmod{2^{4}}$, $97\equiv 1\pmod{2^{5}}$, and $12289\equiv 1\pmod{2^{12}}$ are all prime, we see that $2p^{2}<3^{2^{k-1}}$ whenever $k\geq 4$.
Hence we must have $(k,p)=(2,5)$ or $(3,17)$, and then it is straightforward to see that the only solutions to (5.2) are $k=2$, $c\in\{\pm 2,\pm 3\}$ and $k=3$, $c\in\{\pm 2\}$. In each of those cases, $p$ occurs as $p_{1}$, but then the next prime $\equiv 1\pmod{2^{k}}$ is omitted, as we can see by applying Lemma 4.5 to the equations
$(\pm 10p_{2}\cdots p_{n})^{2}+1=13^{r},\quad(\pm 15p_{2}\cdots p_{n})^{2}+1=2\cdot 13^{r},\quad(\pm 34p_{2}\cdots p_{n})^{4}+1=41^{r}.$